# OBsmith: LLM-powered JavaScript Obfuscator Testing

### v1

@@ 4.1.3. JavaScript Engine
We use Node.js to run our experiments. At its core, Node.js is a server-side JavaScript runtime environment built on the foundation of the V8 JavaScript engine. V8 is developed by Google and used in its Chrome browser, which is responsible for taking the JavaScript code written in a Node.js application and compiling it into the machine code that the computer can execute directly. This compilation process, leveraging V8’s Just-In-Time (JIT) compilation and optimization techniques, allows Node.js applications to achieve high levels of performance and efficiency, particularly in handling concurrent connections and asynchronous operations, says GeeksforGeeks. Essentially, V8 provides the raw power of JavaScript execution, while Node.js extends that power with a rich set of APIs and modules, making it a complete environment for building scalable server-side applications
@@ 4.2. RQ1: LLMs’ Sketch Generation Ability
Answer to RQ1: LLMs’ Sketch Generation Ability GPT-5, Claude 4, and Gemini 2.5 Pro generate sketches that are both syntactically valid and DSL-conformant in the vast majority of cases, with residual issues concentrated in operator specification that our tolerance pass can absorb. Qwen 3 and Grok 4 struggle with domain-specific constraints, producing unrecoverable DSL violations. Providing task context via a code agent improves adherence without changing the underlying LLM.
@@ 4.3.1. Bugs in JS-Confuser
Answer to RQ2: Case Study These findings demonstrate that correctness bugs are pervasive in widely used obfuscators. While some errors manifest as program crashes, others are more insidious—such as silent changes in return values or suppressed exceptions—that can undermine program reliability and complicate debugging. OBsmith’s LLM-driven, sketch-based differential testing proves highly effective at surfacing such issues, reinforcing the need for rigorous correctness validation in obfuscator development.
@@ 4.4. RQ3: Comparison Experiment
Table 2. Comparison with baseline and ablation study
Bug Comparison with baselines Ablation study
OBsmith FuzzJIT Jsfunfuzz Superion DIE Fuzzilli -l -f -e -m -d
constructor name
eval crash
silent fix
error handling
scope violation
control flow change
undefined handling
async crash
class crash
! calss crash
type errors
@@ Results.
Answer to RQ3: Comparison Experiment Under an equal program budget (3,000 programs), OBsmith surfaced all observed correctness bugs, whereas state-of-the-art JavaScript fuzzers did not expose these issues in our setting. The evidence supports using OBsmith alongside engine fuzzers: the former targets transformation-induced semantic drift; the latter remains essential for VM/JIT vulnerabilities.
@@ 4.5. RQ4: Ablation Study
We evaluate how each OBsmith component contributes to bug discovery by enabling exactly one sketch-generation technique or exactly one oracle at a time: the initial LLM sketcher (-l), the feedback loop that refines sketches (-f), automatic sketch extraction from real code (-e), metamorphic testing (-m), and differential testing (-d). Every variant is run under the same program budget as the full system. For the feedback setting, we report only new bugs found by the refined sketches—if the input already exhibits “bug ii” and the loop merely reconfirms bug $i$, we do not count it.
@@ Results.
  • Differential testing (-d) is the single most impactful oracle. It uncovers the majority of classes tied to semantic divergence between the original and obfuscated program—e.g., eval crashes, incorrect undefined handling, type errors, and general error-handling inconsistencies. These require cross-version comparison and are rarely surfaced by engine fuzzers.
@@
Answer to RQ4: Ablation Study (i) Every component except -m contributes to at least one bug class, with -d providing the largest unique lift. (ii) -l, -f, and -e provide complementary coverage: LLM-only sketches (-l) already stress obfuscators; feedback (-f) recovers additional crash types (notably class); and extraction (-e) adds realistic patterns that trigger control/data-flow issues and reveal async crash. (iii) The negative result for -m suggests that general compiler MRs are ill-suited to obfuscator correctness; domain-specific MRs are likely required to make metamorphic testing effective. However, it specifically surfaces brittleness like order-sensitivity and bad interactions between transformation passes that ordinary differential tests can miss. It also complements the differential oracle as a lightweight regression harness across versions/configurations, broadening coverage without needing new ground-truth outputs.
@@ 5. Discussion
Based on investigation from Google(Jiang et al., 2025) and WeChat(Li et al., 2025), runtime performance is also important in JavaScript obfuscation, so beyond correctness of obfuscators, we also evaluate the obfuscation effect on storage runtime performance. To evaluate OBsmith’s ability to surface these trade-offs, we measured average file size, runtime execution time, and memory usage across multiple obfuscation configurations. Table 3 summarizes these results.
@@ 7.1. JavaScript Obfuscator
Previous studies have explored various dimensions of software obfuscation, encompassing obfuscation techniques, detection methodologies, and evaluation frameworks. Sun et al. (Sun et al., 2023) highlighted the significant adverse impact of obfuscation on anti-malware tools. Similarly, Hammad et al. (Hammad et al., 2018) conducted extensive experiments to evaluate the efficacy of leading anti-malware products against diverse obfuscation methods, including seven prominent open-source, academic, and commercial obfuscation tools. Furthermore, various approaches have been proposed specifically to detect JavaScript obfuscation.
However, existing research largely overlooks critical aspects such as evaluating whether obfuscation tools inadvertently alter the original code functionality or negatively impact runtime performance. Skolka et al.(Skolka et al., 2019) studying minified and obfuscated code in the web, however, it study some obsolete test cases and JavaScript TypeScript changes a lot in recent years the thoughts are unavailable now. E.g. Based on Google’s investigation(Jiang et al., 2025), Obfuscator.IO is the most frequently used obfuscator now  90% of obfuscated code was obfuscated by this, but Skolka’s work didn’t cover this work in total. So it’s time to rethink this.
@@
OBsmith uses sketching to catch the domain knowledge. Program sketching (Solar-Lezama et al., 2006, 2007; Solar-Lezama, 2008; Jeon et al., 2015; Solar-Lezama et al., 2008; Singh and Solar-Lezama, 2011; Alur et al., 2013; Singh and Gulwani, 2015; Hong et al., 2025), pioneered by the Sketch system (Solar-Lezama, 2008), offered an exciting new advance in scaling program synthesis, where a partial implementation is given and the goal is to complete it (Bodík and Jobstmann, 2013; Galenson et al., 2014; Feser et al., 2015; Kuncak et al., 2010; Osera and Zdancewic, 2015; Gvero et al., 2011; Mandelin et al., 2005; Feng et al., 2017; Kneuss et al., 2013; Singh and Gulwani, 2015). EdSketch (Hua and Khurshid, 2017) and EdSynth (Yang et al., 2018) defined an optimized backtracking search for completing Java sketches where test executions guided search pruning. The Sketch system provided Java-like Sketch language for writing partial programs, and deployed SAT and inductive synthesis in a counterexample-guided loop to complete them. JSketch enabled sketching Java programs (Jeon et al., 2015) by translating Java to Sketch. PSketch focused on concurrent data structures and enabled sketching them (Solar-Lezama et al., 2008).
@@ 7.3. Large Language Models
LLMs have recently demonstrated strong capabilities in diverse code-related tasks, and they have enabled the automation of many software development and verification themes, including writing code (Bairi et al., 2024; Xia et al., 2025; Zhong et al., 2025a, b), clarifying requirements (Mu et al., 2024), software maintenance (Dilhara et al., 2024; Jin and Lin, 2024; Jiang et al., 2024a; Ma et al., 2024; Xu et al., 2024; Nam et al., 2024), software testing (Ryan et al., 2024; Liu et al., 2024a; Dinella et al., 2022; Jiang et al., 2024b), debugging (Wadhwa et al., 2024; Kang et al., 2024), constructing proofs of theorems in automated provers (First et al., 2023), and human-centric studies (Wang et al., 2024; Imran et al., 2024).
Trained on extensive corpora of code and textual data, these models develop an implicit understanding of programming language syntax, semantics, and code structures(Liu et al., 2024b; Richards et al., 2024; Jiang et al., 2025). Consequently, LLMs exhibit notable capabilities in tasks requiring both natural language comprehension and code manipulation(Jiang et al., 2024b). While models such as StarCoder(Lozhkov et al., 2024) and Gemini demonstrate proficiency in code generation, they remain susceptible to hallucination, wherein they generate plausible but incorrect or nonsensical outputs(Gambardella et al., 2024). Moreover, LLMs exhibit limitations in tasks demanding precise logical and mathematical reasoning(Zhao et al., 2024; Yuan et al., 2023).
@@ 8. Conclusion
In this paper, we present OBsmith, a framework that enables LLM-powered JavaScript obfuscator testing. Using OBsmith, obfuscator developers can use LLMs to generate representative JavaScript sketches which are suitable for testing obfuscators. OBsmith fills sketches and enhances them to enable differential testing. OBsmith also expose program generator as a script to allow developers write their sketches in the same language as the obfuscators they are testing (JavaScript), enabling them to leverage their domain knowledge to set up a code structure likely to lead to obfuscation problem while leaving stakeholders representing expressions they want explored. Using 600 sketches generated by 6 LLMs, OBsmith generates 3,000 programs and found 11 bugs in 2 popular JavaScript obfuscators used by malware developers. These bugs cover obfuscator crash, obfuscated code crash, functionality changes, etc. OBsmith combines the power of LLM to synthesize representative sketches and perform random differential testing to detect critical vulnerabilities.
@@ 9. Data-Availability Statement
We will provide the artifact with all prompts, sketches, programs, and bug report. We will also provide a ready to use artifact to run our sketch filling algorithm, enhance program algorithm, and testing scripts.

### v2

@@ 4.1.3. JavaScript Engine
We use Node.js to run our experiments. At its core, Node.js is a server-side JavaScript runtime environment built on the foundation of the V8 JavaScript engine. V8 is developed by Google and used in its Chrome browser, which is responsible for taking the JavaScript code written in a Node.js application and compiling it into the machine code that the computer can execute directly. This compilation process, leveraging V8’s Just-In-Time (JIT) compilation and optimization techniques, allows Node.js applications to achieve high levels of performance and efficiency, particularly in handling concurrent connections and asynchronous operations. Essentially, V8 provides the raw power of JavaScript execution, while Node.js extends that power with a rich set of APIs and modules, making it a complete environment for building scalable server-side applications.
@@ 4.1.5. Baseline
4.1.6. Test budget.
We budget testing effort by the number of generated candidate programs (i.e., the number of inputs submitted to the obfuscator), rather than by wall-clock time. This choice is motivated by the high variability of LLM generation latency, which can differ substantially across runs and prompts and becomes even less predictable when enabling thinking-style inference. A time budget would therefore conflate testing effectiveness with model-side latency and infrastructure effects, making comparisons between LLM-based generation and traditional fuzzers inherently unfair. Moreover, the two approaches differ fundamentally in their search strategy and computational profile: LLM-based synthesis is driven by token-level inference (sensitive to prompt length, retained context, and inference mode), whereas traditional fuzzers typically explore the input space via lightweight mutations guided by coverage feedback and are primarily bounded by execution throughput. Under a wall-clock budget, these mismatched dynamics would bias results toward whichever approach happens to be cheaper under a particular model and hardware configuration. By contrast, a test-count budget isolates effectiveness per generated input and yields more reproducible comparisons.
@@ 4.2. RQ1: LLMs’ Sketch Generation Ability
Answer to RQ1: GPT-5, Claude 4, and Gemini 2.5 Pro generate sketches that are both syntactically valid and DSL-conformant in the vast majority of cases, with residual issues concentrated in operator specification that our tolerance pass can absorb. Qwen 3 and Grok 4 struggle with domain-specific constraints, producing unrecoverable DSL violations. Providing task context via a code agent improves adherence without changing the underlying LLM.
@@ 4.3.1. Bugs in JS-Confuser
4.3.3. Error-prone obfuscation techniques.
Our failures exhibit clear configuration dependence, but the dominant risk factors are as follows. For JS-Confuser, we observed that more aggressive presets (e.g., Medium/High) are more likely to trigger semantic deviations because they enable transformations that interact poorly with JavaScript’s dynamic semantics and reflective features. A recurring error-prone family concerns exception and unresolvable-identifier handling: obfuscation may replace a required ReferenceError with the value undefined, which can change control flow (e.g., altering which try/catch paths execute) or silently propagate undefined into program state. A second family involves renaming under reflection: aggressive renaming can break eval-based code when string-evaluated expressions refer to original identifiers, and it can change observable identifier metadata such as constructor.name. These observations suggest that JS-Confuser configurations enabling aggressive renaming and exception-related rewrites are comparatively riskier, while less aggressive presets are safer when applications rely on precise exception behavior or reflection.
For Obfuscator.IO, the main configuration-dependent issue we observed is associated with control-flow flattening, which can lead to a runtime TypeError. In contrast, the other crashes we encountered are parsing-stage failures that persist across all tested presets, suggesting that they are less sensitive to configuration choices. We emphasize that these findings are empirical observations from our test suite rather than universal guarantees.
Answer to RQ2: These findings demonstrate that correctness bugs are pervasive in widely used obfuscators. While some errors manifest as program crashes, others are more insidious—such as silent changes in return values or suppressed exceptions—that can undermine program reliability and complicate debugging. OBsmith’s LLM-driven, sketch-based testing proves highly effective at surfacing such issues, reinforcing the need for rigorous correctness validation in obfuscator development.
@@ 4.4. RQ3: Comparison Experiment
Table 5. Comparison with baseline and ablation study, Issue IDs refer to GitHub issues in OWNER/REPO.
Bug [Issue ID] Comparison with baselines Ablation study
OBsmith FuzzJIT Jsfunfuzz Superion DIE Fuzzilli -l -f -e -m -r
constructor name [185]
eval crash [186]
silent fix [188]
error handling [188]
scope violation [189]
control flow change [187]
undefined handling [189]
async crash [1354]
class crash [1354]
! class crash [1354]
type errors [1289]
@@ Results.
Answer to RQ3: Under an equal program budget (3,000 programs), OBsmith surfaced all observed correctness bugs, whereas state-of-the-art JavaScript fuzzers did not expose these issues in our setting. The evidence supports using OBsmith alongside engine fuzzers: the former targets transformation-induced semantic drift; the latter remains essential for VM/JIT vulnerabilities.
@@ 4.5. RQ4: Ablation Study
We evaluate how each OBsmith component contributes to bug discovery by enabling exactly one sketch-generation technique or exactly one oracle at a time: the initial LLM sketcher (-l), the feedback loop that refines sketches (-f), automatic sketch extraction from real code (-e), metamorphic testing (-m), and reference-oriented equivalence testing (-r). Every variant is run under the same program budget as the full system. For the feedback setting, we report only new bugs found by the refined sketches—if the input already exhibits “bug ii” and the loop merely reconfirms bug $i$, we do not count it.
@@ Results.
  • Reference-oriented equivalence testing (-r) is the single most impactful oracle. It uncovers the majority of classes tied to semantic divergence between the original and obfuscated program—e.g., eval crashes, incorrect undefined handling, type errors, and general error-handling inconsistencies. These require cross-version comparison and are rarely surfaced by engine fuzzers.
@@
Answer to RQ4: (1) Every component except -m contributes to at least one bug class, with -r providing the largest unique lift. (2) -l, -f, and -e provide complementary coverage: LLM-only sketches (-l) already stress obfuscators; feedback (-f) recovers additional crash types (notably class); and extraction (-e) adds realistic patterns that trigger control/data-flow issues and reveal async crash. (3) The negative result for -m suggests that general compiler MRs are ill-suited to obfuscator correctness; domain-specific MRs are likely required to make metamorphic testing effective.
@@ 5. Discussion
Based on investigation from Google (Jiang et al., 2026) and WeChat (Li et al., 2025), runtime performance is also important in JavaScript obfuscation, so beyond correctness of obfuscators, we also evaluate the obfuscation effect on storage runtime performance. To evaluate OBsmith’s ability to surface these trade-offs, we measured average file size, runtime execution time, and memory usage across multiple obfuscation configurations. Table 6 summarizes these results.
@@ 7.1. JavaScript Obfuscator
Prior work on software obfuscation spans obfuscation techniques, detection and analysis methods, and empirical evaluation. Sun et al. (Sun et al., 2023) showed that obfuscation can substantially degrade the effectiveness of anti-malware tooling. Hammad et al. (Hammad et al., 2018) performed a broad experimental evaluation of mainstream anti-malware products under a diverse set of obfuscation strategies, covering multiple open-source, academic, and commercial obfuscators. Complementary to these efforts, researchers have also proposed techniques tailored to detecting obfuscation in JavaScript.
However, for JavaScript obfuscators in particular, the literature offers comparatively limited support for evaluating two developer-facing properties: (1) semantic preservation—whether obfuscation changes program behavior—and (2) performance impact—how transformations affect runtime cost. Skolka et al. (Skolka et al., 2019) presented an early measurement study of minified and obfuscated web code, but the JavaScript/TypeScript ecosystem and the set of widely deployed obfuscation tools have changed considerably since then. Google’s recent investigation (Jiang et al., 2026) identifies Obfuscator.IO as a dominant choice and js-confuser is the second choice, suggesting that conclusions drawn from older tool coverage or test cases may not generalize. These observations motivate revisiting evaluation methodology for modern JavaScript obfuscation and developing automated tests that check both behavioral equivalence and performance regressions.
@@
Recent work has explored complementary approaches to improving software reliability, including static analysis, infrastructure for querying heap objects, and compilation-time optimizations (Ding et al., 2026; Thimmaiah et al., 2025; Al Awar et al., 2025; Yuan et al., 2024). Other efforts develop fuzzers that generate test programs directly (Jiang et al., 2023a; Wang et al., 2023b; Xia et al., 2024; Jiang et al., 2023b). In contrast, OBsmith uses sketching to encode domain knowledge. Program sketching (Solar-Lezama et al., 2006, 2007; Solar-Lezama, 2008; Jeon et al., 2015; Solar-Lezama et al., 2008; Singh and Solar-Lezama, 2011; Alur et al., 2013; Singh and Gulwani, 2015; Hong et al., 2025), pioneered by the Sketch system (Solar-Lezama, 2008), offered an exciting new advance in scaling program synthesis, where a partial implementation is given and the goal is to complete it (Bodík and Jobstmann, 2013; Galenson et al., 2014; Feser et al., 2015; Kuncak et al., 2010; Osera and Zdancewic, 2015; Gvero et al., 2011; Mandelin et al., 2005; Feng et al., 2017; Kneuss et al., 2013; Singh and Gulwani, 2015). EdSketch (Hua and Khurshid, 2017) and EdSynth (Yang et al., 2018) defined an optimized backtracking search for completing Java sketches where test executions guided search pruning. The Sketch system provided Java-like Sketch language for writing partial programs, and deployed SAT and inductive synthesis in a counterexample-guided loop to complete them. JSketch enabled sketching Java programs (Jeon et al., 2015) by translating Java to Sketch. PSketch focused on concurrent data structures and enabled sketching them (Solar-Lezama et al., 2008).
@@ 7.3. Large Language Models
LLMs have recently demonstrated strong capabilities in diverse code-related tasks, and they have enabled the automation of many software development and verification themes, including writing code (Bairi et al., 2024; Xia et al., 2025; Zhong et al., 2025a, b), clarifying requirements (Mu et al., 2024), software maintenance (Dilhara et al., 2024; Jin and Lin, 2024; Jiang et al., 2024a; Ma et al., 2024; Xu et al., 2024; Nam et al., 2024), software testing (Ryan et al., 2024; Liu et al., 2024; Dinella et al., 2022; Jiang et al., 2024b; Ni and Li, 2025), debugging (Wadhwa et al., 2024; Kang et al., 2024), constructing proofs of theorems in automated provers (First et al., 2023), and human-centric studies (Wang et al., 2024; Imran et al., 2024).
Trained on extensive corpora of code and textual data, these models develop an implicit understanding of programming language syntax, semantics, and code structures (Liu et al., 2025; Jiang et al., 2026). Consequently, LLMs exhibit notable capabilities in tasks requiring both natural language comprehension and code manipulation (Jiang et al., 2024b). While models such as StarCoder (Lozhkov et al., 2024) and Gemini demonstrate proficiency in code generation, they remain susceptible to hallucination, wherein they generate plausible but incorrect or nonsensical outputs (Gambardella et al., 2024). Moreover, LLMs exhibit limitations in tasks demanding precise logical and mathematical reasoning (Zhao et al., 2024; Yuan et al., 2023).
@@ 8. Conclusion
In this paper, we present OBsmith, a framework that enables LLM-powered JavaScript obfuscator testing. Using OBsmith, obfuscator developers can use LLMs to generate representative JavaScript sketches which are suitable for testing obfuscators. OBsmith fills sketches and enhances them to enable reference-oriented equivalence testing. OBsmith also exposes program generator as a script to allow developers to write their sketches in the same language as the obfuscators they are testing (JavaScript), enabling them to leverage their domain knowledge to set up a code structure likely to lead to obfuscation problems. Using 600 sketches generated by 6 LLMs, OBsmith generates 3,000 programs and found 11 bugs in 2 popular JavaScript obfuscators used by malware developers. These bugs cover obfuscator crash, obfuscated code crash, control flow changes, etc.
@@ 9. Data-Availability Statement
The sketch filling algorithm, enhance algorithm, testing scripts, prompts, sketches, and generated programs for OBsmith are publicly available at https://github.com/shanjiang98/obsmith.
Acknowledgements.
We would like to acknowledge Department of Energy award DE-SC0024467 for their support.
Appendix A Appendix
<context> You are a software testing engineer. Now your task is to generate some template JavaScript programs, and these programs will be used to test JavaScript obfuscators. In template, you should use holes to represent a holder, these holes can be numberLiteral, or booleanLiteral. And these holes can also be numberReference, or booleanReference. And you can use binary expression including arithmetic operation(e.g. +, -, *, /, ...), relation operation(>, <, ==, ===, >=, <=, ...) and logic operation (AND, OR, NOT, ...) in templates, the operands can be any type of literals, references, or holes. I will use templates you generate to do differential testing. Specifically, I will fill holes to get programs and run the filled programs with JavaScript obfuscators and minifiers. Finally I will compare if the transformed JavaScript program have the same output of original filled program. </context> <example> <template> let s1 = numberLiteral; let s2 = numberLiteral; let b1 = booleanLiteral; let arr = [s1++, s2, numberLiteral, numberLiteral, booleanLiteral] for (int i = 0; i < arr.length; i++) { if (logic(relation(numberReference, numberReference, <=), relation(numberReference, numberReference, <=), &&, ||)){ arr[i] = arithmetic(numberReference, numberReference, +, *); } if (booleanReference) { booleanReference = booleanLiteral; booleanReference = booleanReference; } } </template> <filled program> let s1 = 1098733; let s2 = 387923l2; let b1 = true; let arr = [s1++, s2, 37289, 776, true] for (int i = 0; i < arr.length; i++) { if (arr[1] <= s2 || s2 <= arr[0]){ arr[i] = arr[0] + s1; } if (arr[4]){ b1 = false; b1 = arr[4]; } } </filled program> </examples> <format> You should output sketches in a markdown block without any explaination. e.g.: ‘‘‘javascript <sketch 1> ... <sketch 10> ‘‘‘ </format> <instruction> Now please generate <number> different templates for me. You should use templates to cover as many representative javascript programs as possible, including different control flows, common functions, etc. Make sure your output follows format instruction. </instruction> Figure 5. OBsmith prompt for LLM-based sketch generation